Determine the pH of the following aqueous solution
Answer #1

12 )

a )

pKa = -logKa = 3.34

pH = 1/2 [pKa -logC]

pH = 1/2 [3.34 - log 0.250]

pH = 1.97

b)

pOH = 1/2 [pKb - log C]

= 1/2 [4.74 -log 0.175]

= 2.75

pH + pOH = 14

pH = 11.25